OpenStudy (anonymous):

how can i find the sum or difference of 22-2n/2?

OpenStudy (aum):

The question is not clear. Could you type the full problem as it appears?

OpenStudy (aum):

To find the difference you need two terms. What are the two terms?

OpenStudy (anonymous):

22-2n under 4

OpenStudy (anonymous):

exa

OpenStudy (anonymous):

the two terms is 22 and 2 i think :/

OpenStudy (aum):

First term is \[\large \frac{4}{22-2n} \]?

OpenStudy (aum):

You typed 22-2n under 4. Does it look like the one above?

OpenStudy (anonymous):

oops yes that was a accident my previous message its under 2

OpenStudy (aum):

\[\large \frac{22-2n}{2} = 11 - n \]

OpenStudy (anonymous):

is this the answer?

OpenStudy (aum):

All I did what simplify what you gave. I did not find the sum or difference. Just simplified it. Since the question is not clear IDK what the answer is.

OpenStudy (anonymous):

what do you mean? my question was very clear when i drew the problem. my textbook said to find the sum or difference so that's why i asked this question.

OpenStudy (anonymous):

I'm sorry if my question wasn't very clear.

OpenStudy (anonymous):

what about write a fraction as a sum or difference?

OpenStudy (anonymous):

are you still here?

OpenStudy (aum):

Normally you are given two quantities which you can add together and find the sum. Or you can subtract one quantity from another and find the difference. Like find the sum of 5 and 2. It is 5 + 2 = 7 or 5 - 2 = 3 But to give something like (5-2)/3 and say find the sum or difference is not really clear.

OpenStudy (aum):

\[ \large \frac{22-2n}{2} = \frac{22}{2} - \frac{2n}{2} = 11 - n \]is the best I can do.